OpenStudy (anonymous):

The number of right triangles with sides of lengths \[\sqrt{n+1}\], \[\sqrt{n+2}\], and \[\sqrt{n+3}\] is: a) 0 b) 1 c) 2 d) 3 e) unbounded

hartnn (hartnn):

\(\sqrt{n+3}\) is the longest side, so its the hypotenuse. now apply pythagoras theorem.

OpenStudy (anonymous):

\[\sqrt{n+1}^{2}+\sqrt{n+2}^{2}=\sqrt{n+3}^{2}\]

OpenStudy (anonymous):

(n+1)+(n+2)=(n+3)

hartnn (hartnn):

yeah, how many values of 'n' satisfy that equation ?

OpenStudy (anonymous):

2n+3=n+3 n=0

hartnn (hartnn):

number of right triangles = number of values of n

OpenStudy (anonymous):

0 right triangles?

hartnn (hartnn):

no, n=0 is one value of n. so there's one right triangle with side, sqrt 1 sqrt 2 and sqrt 3

OpenStudy (anonymous):

oh, so there is 1 right triangle, where n = 0

hartnn (hartnn):

yup.

OpenStudy (anonymous):

Thank you! :)